# AsEP: Benchmarking Deep Learning Methods for Antibody-specific Epitope Prediction

### v1

@@ Abstract
Epitope identification is vital for antibody design yet challenging due to the inherent variability in antibodies. While many deep learning methods have been developed for general protein binding site prediction tasks, whether they work for epitope prediction remains an understudied research question. The challenge is also heightened by the lack of a consistent evaluation pipeline with sufficient dataset size and epitope diversity. We introduce a filtered antibody-antigen complex structure dataset, AsEP (Antibody-specific Epitope Prediction). AsEP is the largest of its kind and provides clustered epitope groups, allowing the community to develop and test novel epitope prediction methods. AsEP comes with an easy-to-use interface in Python and pre-built graph representations of each antibody-antigen complex while also supporting customizable embedding methods. Based on this new dataset, we benchmarked various representative general protein-binding site prediction methods and find that their performances are not satisfactory as expected for epitope prediction. We thus propose a new method, WALLE, that leverages both protein language models and graph neural networks. WALLE demonstrate about $5$X performance gain over existing methods. Our empirical findings evidence that epitope prediction benefits from combining sequential embeddings provided by language models and geometrical information from graph representations, providing a guideline for future method design. In addition, we reformulate the task as bipartite link prediction, allowing easy model performance attribution and interpretability. We open-source our data and code at https://github.com/biochunan/AsEP-dataset.
@@ 1 Introduction
Antibodies are specialized proteins produced by our immune system to combat foreign substances called antigens. Their unique ability to bind with high affinity and specificity sets them apart from regular proteins and small-molecule drugs, making them increasingly popular in therapeutic engineering.
While the community is shifting toward computational antibody design given a pre-determined epitope  (Jin et al., 2022; Zhou et al., 2024; Bennett et al., 2024), accurate prediction of epitopes remains underexplored. Accurate identification of epitopes is beneficial for understanding antibody-antigen interactions, antibody function, and streamlining antibody engineering. This task remains challenging due to multiple factors (Akbar et al., 2022; Hummer et al., 2022), such as, the lack of comprehensive datasets, limited interpretability and generalizability. Available datasets are limited in size, up to 582 complexes from Bepipred-3.0 (Clifford et al., 2022), and feature disproportionate representation among various epitopes.
Existing methods perform poorly on epitope prediction task (Cia et al., 2023), with a ceiling MCC (Matthew’s Correlation Coefficient) of $0.06$. However, recent advancements in graph-based learning, coupled with an increase in available antibody structures in the Protein Data Bank (PDB) (Berman et al., 2003), highlight the need to reevaluate current methods and establish a benchmark dataset for predicting antibody-antigen interactions.
We approach the problem as a bipartite graph link prediction task. Traditionally, graph link prediction focuses on identifying connections within the same graph, such as in protein-protein interaction networks. Our research extends this concept to bipartite graphs at the molecular level and proposes our own model, WALLE. Because existing methods generally predict protein binding sites or epitopes in antibody-antigen complexes rather than residue-residue interactions, we focus on benchmarking the node classification task while providing WALLE’s performance on the bipartite link prediction task as a baseline.
@@ 2 Previous Work
Accurate epitope prediction for antibody design remains challenging due to the complexity of antibody-antigen interactions and limitations of existing datasets and methods. Several computational approaches have been developed, but they often fall short in terms of accuracy and applicability. Here, we present a representative set of state-of-the-art methods.
EpiPred (Krawczyk et al., 2014) implements a graph-based antibody-antigen specific scoring function that considers all possible residue-residue pairs at the interface between anybody and antigen structures. It samples surface patches from the antigen structure surface and selects the highest-ranked patch with the highest score as the predicted set of epitope residues.
ESMFold(Lin et al., 2023) is a protein language model based on ESM2 (Lin et al., 2023), and its folding head was trained on over 325 thousand protein structures. It achieves comparable performance to AlphaFold2 (Jumper et al., 2021) and is included in our benchmarking due to its faster processing.
@@
PECAN and EPMP (Pittala & Bailey-Kellogg, 2020; Vecchio et al., 2021) are graph neural networks that predict epitope residues taken antibody-antigen structure pairs as input. They use position-specific scoring matrices (PSSM) as node embeddings and graph attention networks to predict the binary labels of nodes in the antigen graph. For details, we refer readers to  Section A.1.
@@
Zhao et al. (2024) benchmarked docking methods like ZDOCK (Pierce et al., 2011), ClusPro (Kozakov et al., 2017), and HDOCK (Yan et al., 2017), and Alphafold-Multimer (Evans et al., 2022) on a set of $112$ antibody-antigen complexes. They showed that all docking methods gave a success rate of $8.0\%$ at most if using the top 5 decoys; Alphafold-Multimer showed a better performance with a $15.3\%$ success rate.
@@ 3 Problem Formulation
Antibody-antigen interaction is important for analyzing protein structures. The problem can be formulated as a bipartite graph link prediction task. The inputs are two disjoint graphs, an antibody graph $G_{A}=\left(V_{A},E_{A}\right)$ and an antigen graph $G_{B}=\left(V_{B},E_{B}\right)$, where $V_{x}$ is the vertice set for graph $x$ and $E_{x}$ is the edge set for graph $x$. Since the neural networks only take continuous values as input, we also have a function $h$ to encode each vertex into a vector $h:V\to\mathbb{R}^{D}$. The design of the encoding function depends on the methods. For example, $h$ can be a one-hot encoding layer or pretrained embeddings given by a protein language model. We use different encoding functions for antibodies and antigens: $h_{A}:V_{A}\to\mathbb{R}^{D_{A}}$, and $h_{B}:V_{B}\to\mathbb{R}^{D_{B}}$.
In addition, $E_{A}\in\{0,1\}^{|V_{A}|\times|V_{A}|}$ and $E_{B}\in\{0,1\}^{|V_{B}|\times|V_{B}|}$ denote the adjacency matrices for the antibody and antigen graphs, respectively. In this work, the adjacency matrices are calculated based on the distance matrix of the residues. Each entry $e_{ij}$ denotes the proximity between residue $i$ and residue $j$; $e_{ij}=1$ if the Euclidean distance between any non-hydrogen atoms of residue $i$ and residue $j$ is less than $4.5$Å, and $e_{ij}=0$ otherwise. The antibody graph $G_{A}$ is constructed by combining the CDR residues from the heavy and light chains of the antibody, and the antigen graph $G_{B}$ is constructed by combining the surface residues of the antigen. The antibody and antigen graphs are disjoint, i.e., $V_{A}\cap V_{B}=\emptyset$.
Figure 1: An example to illustrate interacting residues. The two dashed lines denote the distance between non-hydrogen atoms from different interacting residues from two different protein chains.
@@
This classification takes into account the structure of the antibody graph, $G_{A}$, mirroring the specificity of antibody-antigen binding interactions. Different antibodies can bind to various antigen locations, corresponding to varying subsets of epitope nodes in $G_{B}$. This question differs from conventional epitope prediction methods that do not consider the antibody structure and end up predicting the likelihood of the subset of antigen nodes serving as epitopes, such as ScanNet (Tubiana et al., 2022), MaSIF (Gainza et al., 2020).
The task is to develop a binary classifier $f:V_{B}\to\{0,1\}$ that takes both the antibody and antigen graphs as input and predicts the label for antigen nodes as formulated below:
@@
Bipartite Link Prediction The second task takes it further by predicting concrete interactions between nodes in $G_{A}$ and $G_{B}$, resulting in a bipartite graph that represents these antibody-antigen interactions. Moreover, this helps attribute the model performance to specific interactions at the molecular level and provide more interpretability. Accurately predicting these interactions is critical for understanding the binding mechanisms and for guiding antibody engineering.
We model the antibody-antigen interaction as a bipartite graph:
$K_{m,n}=(V_{A},V_{B},E)$
where $m=|V_{A}|$ and $n=|V_{B}|$ denote the number of nodes in the two graphs, respectively, and $E$ denotes all possible inter-graph links. In this bipartite graph, a node from the antibody graph is connected to each node in the antigen graph via an edge $e\in E$. The task is then to predict the label of each bipartite edge. If the residues of a pair of nodes are located within $4.5$Å of each other, referred to as in contact, the edge is labeled as $1$; otherwise, 00. For any pair of nodes, denoted as $(v_{a},v_{b})\ \forall v_{a}\in V_{A},v_{b}\in V_{B}$, the binary classifier $g:K_{m,n}\rightarrow\{0,1\}$ is formulated as below:
@@ 4.3 Dataset split
We propose two types of dataset split settings. The first is a random split based on the ratio of epitope to antigen surface residues, #epitope nodes#antigen $\frac{\#\text{epitope nodes}}{\#\text{antigen nodes}}$; the second is a more challenging setting where we split the dataset by epitope groups. The first setting is straightforward and used by previous methods. The second setting is more challenging because it requires the model to generalize to unseen epitope groups.
Split by epitope to antigen surface ratio As aforementioned, the number of non-interface nodes in the antigen graph is much larger than the number of interface nodes. While epitopes usually have a limited number of residues, typically around $14.6\pm 4.9$ amino acids (Reis et al., 2022), the antigen surface may extend to several hundred or more residues. The complexity of the classification task therefore increases with the antigen surface size. To ensure similar complexity among train, validation, and test sets, we stratified the dataset to include a similar distribution of epitope to non-epitope nodes in each set. Table S3 shows the distribution of epitope-to-antigen surface ratios in each set. This led to $1383$ antibody-antigen complexes for the training set and $170$ complexes each for the validation and test sets. The list of complexes in each set is provided in the Supplementary Table SI-split-epitope-ratio.csv.
@@
We inlcude $641$ unique antigens and $973$ epitope groups in our dataset. We include multi-epitope antigens. For example, there are $64$ distinct antibodies that bind to coronavirus spike protein.
We can see that different antibodies bind to different locations on the same antigen. Details of all epitope groups are provided in the Supplementary Table SI-AsEP-entries.csv. We then split the dataset into train, validation, and test sets such that the epitopes in the test set are not found in either train or validation sets. We followed an 80%/10%/10% split for the number of complexes in each set. This resulted in $1383$ complexes for the training set and $170$ complexes for the validation and test sets. The list of complexes in each set is provided in the Supplementary Table SI-split-epitope-group.csv.
@@
5 WALLE: a graph-based method for epitope prediction
Alongside our dataset interface, we also provide a graph-based model named WALLE. It takes as input a pair of antibody and antigen graphs, constructed as detailed above for the AsEP dataset, and makes node- and edge-level predictions.
@@ 5 WALLE: a graph-based method for epitope prediction
Graph Modules The architecture of WALLE incorporates graph modules that process the input graphs of antibody and antigen structures, as depicted in Figure 3). Inspired by PECAN and EPMP, our model treats the antibody and antigen graphs separately, with distinct pathways for each. The antibody graph is represented by node embeddings $X_{A}$ with a shape of $(M,512)$ and an adjacency matrix $E_{A}$, while the antigen graph is described by node embeddings $X_{B}$ with a shape of $(N,480)$ and its corresponding adjacency matrix $E_{B}$. The embedding sizes are consistent with AntiBERTy and ESM2 (esm2_t12_35M_UR50D).
@@
Both antibody and antigen graph nodes are first projected into the dimensionality of $128$ using fully connected layers. The resulting embeddings are then passed through two Graph Convolutional Network (GCN) modules consecutively to refine the features and yield updated node embeddings $X_{A}^{\prime}$ and $X_{B}^{\prime}$ with a reduced dimensionality of $(M,64)$. The output from the first GCN layer is passed through a ReLU activate function. Outputs from the second GCN layer are directly fed into the Decoder module. These GCNs operate independently, each with its own parameters, ensuring that the learned representations are specific to the antibody or the antigen.
The use of separate GCN modules for the antibody and antigen allows for the capture of unique structural and functional characteristics pertinent to each molecule before any interaction analysis. This design choice aligns with the understanding that the antibody and antigen have distinct roles in their interactions and their molecular features should be processed separately.
Decoder
@@
We used a simple decoder to predict the binary labels of edges between the antibody and antigen graphs, which takes a pair of node embeddings output by the graph modules as input and predicts the probability of each edge. An edge is assigned a binary label of $1$ if the predicted probability is greater than $0.5$ or 00 otherwise. This is shown as the Decoder module in Figure 3. For the epitope prediction task, we convert edge-level predictions to node-level by summing the predicted probabilities of all edges connected to an antigen node; we assign the antigen node a label of $1$ if the number of connected edges is greater than a threshold or 00 otherwise. The threshold is treated as a hyperparameter and is optimized in the experiments.
Implementation
@@
We used PyTorch Geometric (Fey & Lenssen, 2019) framework to build our model. The graph modules are implemented using the GCNConv module from PyTorch Geometric. We trained the model to minimize a loss function consisting of two parts: a weighted binary cross-entropy loss for the bipartite graph link reconstruction and a regularizer for the number of positive edges in the reconstructed bipartite graph. We used the same set of hyperparameters and loss functions for both dataset settings. The loss function and hyperparameters are described in detail in Section A.6.

### v2

@@ Abstract
Epitope identification is vital for antibody design yet challenging due to the inherent variability in antibodies. While many deep learning methods have been developed for general protein binding site prediction tasks, whether they work for epitope prediction remains an understudied research question. The challenge is also heightened by the lack of a consistent evaluation pipeline with sufficient dataset size and epitope diversity. We introduce a filtered antibody-antigen complex structure dataset, AsEP (Antibody-specific Epitope Prediction). AsEP is the largest of its kind and provides clustered epitope groups, allowing the community to develop and test novel epitope prediction methods and evaluate their generalisability. AsEP comes with an easy-to-use interface in Python and pre-built graph representations of each antibody-antigen complex while also supporting customizable embedding methods. Using this new dataset, we benchmark several representative general protein-binding site prediction methods and find that their performances fall short of expectations for epitope prediction. To address this, we propose a novel method, WALLE, which leverages both unstructured modeling from protein language models and structural modeling from graph neural networks. WALLE demonstrate up to $3$-$10$X performance improvement over the baseline methods. Our empirical findings suggest that epitope prediction benefits from combining sequential features provided by language models with geometrical information from graph representations. This provides a guideline for future epitope prediction method design. In addition, we reformulate the task as bipartite link prediction, allowing convenient model performance attribution and interpretability. We open source our data and code at https://github.com/biochunan/AsEP-dataset.
@@ 1 Introduction
Antibodies are specialized proteins produced by our immune system to combat foreign substances called antigens. Their unique ability to bind with high affinity and specificity sets them apart from regular proteins and small-molecule drugs, making them increasingly popular in therapeutic engineering. While the community is shifting towards computational antibody design based on pre-determined epitopes (Jin et al., 2022; Zhou et al., 2024; Bennett et al., 2024), accurate prediction of epitopes themselves remains underexplored. Precise epitope identification is essential for understanding antibody-antigen interactions and antibody functions, as well as streamlining antibody engineering. The task remains challenging due to multiple factors, including the lack of comprehensive datasets, limited interpretability, and low generalizability (Akbar et al., 2022; Hummer et al., 2022). Existing datasets are limited in size, e.g. only $582$ complexes in Bepipred-$3.0$ (Clifford et al., 2022), and often exhibit disproportionate representation among different epitopes. Current methods perform poorly on epitope prediction (Cia et al., 2023), with a ceiling MCC (Matthew’s Correlation Coefficient) of $0.06$. Besides, recent advances in graph learning algorithms, along with an increase in available antibody structures in the Protein Data Bank (PDB) (Berman et al., 2003), highlight the need to reevaluate current methods and establish a new benchmark dataset for predicting antibody-antigen interactions.
We approach the problem as a bipartite graph link prediction task, where the goal is to directly identify connections between two distinct graphs representing the antibody and antigen. Unlike conventional link prediction tasks (Nickel et al., 2016; Trouillon et al., 2016; Zhang & Chen, 2018; Chen et al., 2021), aiming to predict links within an individual graph (potentially with multiple relation types), such as in a protein-protein interaction network, our approach predicts links across a pair of molecular graphs. Our model, WALLE, is designed to predict fine-grained residue-residue interactions, i.e. bipartite links between the antibody and antigen, while it can also adapt to solve the coarser node classification task, distinguishing binding nodes from non-binding ones. Since most existing methods focus on predicting protein binding sites (i.e. node classification), we first benchmark this binding node classification task for straightforward comparisons with these methods. Moreover, we include WALLE’s performance on the bipartite link prediction task as a baseline for future work.
@@ 2 Previous Work
Accurate epitope prediction for antibody design remains challenging due to the complexity of antibody-antigen interactions and limitations of existing datasets and methods. Several computational approaches have been developed, but they often fall short in terms of accuracy and applicability. Here, we present a representative, yet non-exhaustive, set of state-of-the-art methods.
EpiPred (Krawczyk et al., 2014) implements a graph-based antibody-antigen specific scoring function that considers all possible residue-residue pairs at the interface between antibody and antigen structures. It samples surface patches from the antigen structure and selects the highest-ranked patch as the predicted set of epitope residues.
ESMFold(Lin et al., 2023) is a protein language model based on ESM2 (Lin et al., 2023), and its folding head was trained on over 325 thousand protein structures. It achieves comparable performance to AlphaFold2 (Jumper et al., 2021) and is included in our benchmark due to its faster processing.
@@
PECAN and EPMP (Pittala & Bailey-Kellogg, 2020; Vecchio et al., 2021) are graph neural networks that predict epitope residues taking antibody-antigen structure pairs as input. They use position-specific scoring matrices (PSSM) as node embeddings and graph attention networks to predict the binary labels of nodes in the antigen graph. For details, we refer readers to  Section A.1.
@@
Zhao et al. (2024) benchmarked docking methods like ZDOCK (Pierce et al., 2011), ClusPro (Kozakov et al., 2017), and HDOCK (Yan et al., 2017), and Alphafold-Multimer (Evans et al., 2022) on a set of $112$ antibody-antigen complexes. They showed that all docking methods gave a success rate of $8.0\%$ at most if using the top 5 decoys; AlphaFold-Multimer showed a better performance with a $15.3\%$ success rate, so we included AlphaFold-Multimer (version 2.3) but benchmarked on a separate subset of the proposed dataset (AsEP) according to its training data cutoff date.
@@ 3 Problem Formulation
Antibody-antigen interaction is important for analyzing protein structures. The problem can be formulated as a bipartite graph link prediction task. The inputs are two disjoint graphs, an antibody graph $G_{A}=\left(V_{A},E_{A}\right)$ and an antigen graph $G_{B}=\left(V_{B},E_{B}\right)$, where $V_{x}$ is the vertice set for graph $x$ and $E_{x}$ is the edge set for graph $x$. Since neural networks only take continuous values as input, we encode each vertex into a vector with the function $h:V\to\mathbb{R}^{D}$. The design choice of the encoding function depends on the methods. For example, $h$ can be a one-hot encoding layer or pretrained embeddings given by a protein language model. We use different encoding functions for antibodies and antigens: $h_{A}:V_{A}\to\mathbb{R}^{D_{A}}$, and $h_{B}:V_{B}\to\mathbb{R}^{D_{B}}$.
In addition, $E_{A}\in\{0,1\}^{|V_{A}|\times|V_{A}|}$ and $E_{B}\in\{0,1\}^{|V_{B}|\times|V_{B}|}$ denote the adjacency matrices for the antibody and antigen graphs, respectively. In this work, the adjacency matrices are calculated based on the distance matrix of the residues. Each entry $e_{ij}$ denotes the proximity between residue $i$ and residue $j$; $e_{ij}=1$ if the Euclidean distance between any non-hydrogen atoms of residue $i$ and residue $j$ is less than $4.5$Å, and $e_{ij}=0$ otherwise (See example in Figure 1. The antibody graph $G_{A}$ is constructed by combining the CDR residues from the heavy and light chains of the antibody, and the antigen graph $G_{B}$ is constructed by combining the surface residues of the antigen. The antibody and antigen graphs are disjoint, i.e., $V_{A}\cap V_{B}=\emptyset$.
Figure 1: An example illustrating interacting residues. The two dashed lines indicate distances between non-hydrogen atoms from different interacting residues across two protein chains, with each chain’s carbon atoms colored cyan and green.
@@
This classification takes into account the structure of the antibody graph, $G_{A}$, mirroring the specificity of antibody-antigen binding interactions. Different antibodies can bind to various antigen locations, corresponding to varying subsets of epitope nodes in $G_{B}$. This formulation differs from conventional antigen-only epitope prediction that does not consider the antibody structure and ends up predicting the likelihood of the subset of antigen nodes serving as epitopes, such as ScanNet (Tubiana et al., 2022), MaSIF (Gainza et al., 2020). The goal is to develop a binary classifier $f:V_{B}\to\{0,1\}$ that takes both antibody and antigen graphs as input and predicts the labels for antigen nodes:
@@
Bipartite Link Prediction The second task takes it further by predicting concrete interactions between nodes in $G_{A}$ and $G_{B}$, resulting in a bipartite graph that represents these antibody-antigen interactions. Moreover, this helps attribute the model performance to specific interactions at the molecular level and provide more interpretability. Accurately predicting these interactions is critical for understanding the binding mechanisms and for guiding antibody engineering. We model the antibody-antigen interaction as a bipartite graph $K_{m,n}=(V_{A},V_{B},E)$ where $m=|V_{A}|$ and $n=|V_{B}|$ denote the numbers of nodes in the two graphs, respectively, and $E$ denotes all possible inter-graph links. In this bipartite graph, a node from the antibody graph is connected to each node in the antigen graph via an edge $e\in E$. The task is then to predict the label of each bipartite edge. If the residues of a pair of nodes are located within $4.5$Å of each other, referred to as in contact, the edge is labeled as $1$; otherwise, 00. For any pair of nodes, denoted as $(v_{a},v_{b})\ \forall v_{a}\in V_{A},v_{b}\in V_{B}$, the binary classifier $g:K_{m,n}\rightarrow\{0,1\}$ is formulated as below:
@@ 4.3 Dataset split
We propose two types of dataset split settings. The first is a random split based on the ratio of epitope to antigen surface residues, #epitope nodes#antigen $\frac{\#\text{epitope nodes}}{\#\text{antigen nodes}}$; the second is a more challenging setting where we split the dataset by epitope groups. The first setting is straightforward and used by previous methods, while the second setting requires the model to generalize to unseen epitope groups.
Split by epitope to antigen surface ratio As aforementioned, the number of non-interface nodes in the antigen graph is much larger than the number of interface nodes. While epitopes usually have a limited number of residues, typically around $14.6\pm 4.9$ amino acids (Reis et al., 2022), the antigen surface may extend to several hundred or more residues. The complexity of the classification task, therefore, increases with the antigen surface size. To ensure similar complexity among train, validation, and test sets, we stratified the dataset to include a similar distribution of epitope to non-epitope nodes in each set. Table S3 shows the distribution of epitope-to-antigen surface ratios in each set. This led to $1383$ antibody-antigen complexes for the training set and $170$ complexes each for the validation and test sets. The list of complexes in each set is provided in the Supplementary Table SI-split-epitope-ratio.csv.
@@
We inlcude $641$ unique antigens and $973$ epitope groups in our dataset. We include multi-epitope antigens. For example, there are $64$ distinct antibodies that bind to coronavirus spike protein. We can see that different antibodies bind to different locations on the same antigen. Details of all epitope groups are provided in the Supplementary Table SI-AsEP-entries.csv. We then split the dataset into train, validation, and test sets such that the epitopes in the test set are not found in either train or validation sets. We used an 80%/10%/10% split for the number of complexes in each set. This resulted in $1383$ complexes for the training set and $170$ complexes for the validation and test sets. The list of complexes in each set is provided in the Supplementary Table SI-split-epitope-group.csv.
@@
5 WALLE: A Hybrid Method for Epitope Prediction
Alongside our dataset interface, we also provide a new method named WALLE. It takes as input a pair of antibody and antigen graphs, constructed as detailed above for the AsEP dataset, and makes node-level and edge-level predictions.
@@ 5 WALLE: A Hybrid Method for Epitope Prediction
Graph Modules The architecture of WALLE incorporates graph modules that process the input graphs of antibody and antigen structures, as depicted in Figure 3). Inspired by PECAN and EPMP, our model treats the antibody and antigen graphs separately, with distinct pathways for each. The antibody graph is represented by node embeddings $X_{A}$ with a shape of $(M,D_{A})$ and an adjacency matrix $E_{A}$, while the antigen graph is described by node embeddings $X_{B}$ with a shape of $(N,D_{B})$ and its corresponding adjacency matrix $E_{B}$. Both antibody and antigen graph nodes are first projected into the dimensionality of $128$ using fully connected layers. The resulting embeddings are then passed through two GNN layers consecutively to refine the features and yield updated node embeddings $X_{A}^{\prime}$ and $X_{B}^{\prime}$ with a reduced dimensionality of $(M,64)$. The output from the first GNN layer is passed through a ReLU activation function. Outputs from the second GNN layer are directly fed into the Decoder module. These layers operate independently, each with its own parameters, ensuring that the learned representations are specific to the antibody or the antigen. The use of separate graph modules for the antibody and antigen allows for the capture of unique structural and functional characteristics pertinent to each molecule before any interaction analysis. This design choice aligns with the understanding that antibodies and antigens have distinct roles in their interactions and that their molecular features should be processed separately.
Combining unstructured (sequential) with structural modeling
The embedding size, $D_{A}$ and $D_{B}$, are determined by the pre-trained protein language model (PPLM). We extracted embeddings from the final layer outputs of each PPLM as node features for our GNNs. We experimented with different combinations of graph neural network (GNN) architectures and PPLM embeddings. We assessed commonly used graph modules, Graph Convolutional Network (GCN) (Kipf & Welling, 2017), Graph Attention Network (GAT) (Veličković et al., 2018), and GraphSAGE networks (Hamilton et al., 2018); and PPLMs including, AntiBERTy (Ruffolo et al., 2021), ESM2-35M (esm2_t12_35M_UR50D), ESM2-650M (esm2_t33_650M_UR50D) (Lin et al., 2022).
@@
Decoder We used a simple decoder to predict binary labels for edges between the antibody and antigen graphs. This decoder takes pairs of node embeddings, output by the graph modules, as input and predicts the probability of each edge. During hyperparameter tuning, we compute the final logits by either taking the inner product of the antibody and antigen embeddings or concatenating them and passing them through a single linear layer with dropout. A sigmoid function is applied to produce the final probabilities. An edge is assigned a binary label of $1$ if the predicted probability exceeds $0.5$ or 00 otherwise. This is shown as the Decoder module in Figure 3. For the epitope prediction task, we convert edge-level predictions to node-level by summing the predicted probabilities of all edges connected to an antigen node. We assign the antigen node a label of $1$ if the number of connected edges surpasses a set threshold or 00 otherwise. This threshold is treated as a hyperparameter and optimized during experimentation.
@@
Implementation We used PyTorch Geometric (Fey & Lenssen, 2019) framework to build our model. The graph modules are implemented using GCNConv. GATConv, SAGEConv modules from PyTorch Geometric. We trained the model to minimize a loss function consisting of two parts: a weighted binary cross-entropy loss for the bipartite graph link reconstruction and a regularizer for the number of positive edges in the reconstructed bipartite graph. We used the same set of hyperparameters and loss functions for both dataset settings. The loss function and hyperparameters are described in detail in Section A.6. We report only the best performance of different combinations of graph modules and pre-trained language model embeddings in Table 1 (refer to Appendix B for the performance of all combinations.
6 Results and Discussion
For epitope residue prediction, we evaluated each method on both dataset split settings using the metrics described in Section A.2. The results are summarized in Table 1(a) and Table 1(b), showing the average performance metrics across the test set samples. Since various combinations of graph architectures and pre-trained language models were tested for node embeddings, we report only the performance of the best-performing combination in each table. WALLE generally outperforms other methods across all metrics on both dataset splits. As previously noted, existing methods do not evaluate interaction prediction; therefore, the baseline performance for bipartite link prediction is provided in Table S8 for future reference.